### v1

@@ 2 Sample selection
The BLOeM sample (see Fig. 1) was selected from the third Gaia data release catalogue (Gaia DR3, Gaia Collaboration et al., 2023). The catalogue was retrieved from the Gaia database using a search radius of around the SMC center ( = 00:52:38, -72:48:01; epoch J2000). To achieve S/N per pixel (0.2 Å spectral bin), only stars with mag were retrieved. Foreground objects were filtered via two constraints. First, the parallax was required to be consistent with zero111Gaia is not sensitive enough to measure non-vanishing parallaxes within errors at the SMC distance. within 5, that is . Second, the proper motions of the stars were required to be consistent within 15 with the SMC proper motion ( and ) following Yang et al. (2019) and Schootemeijer et al. (2021).
@@
We omitted the 12 known SMC Wolf-Rayet (WR) stars (Neugent et al., 2018) from the sample, since they have been previously monitored (Foellmi et al., 2003; Hainich et al., 2015; Shenar et al., 2016; Schootemeijer et al., 2024). Finally, we omitted potential red supergiants (RSGs) from the sample by imposing mag: due to the large radii of RSGs, RSG binaries have periods that exceed a few years (e.g. Patrick et al., 2019; Neugent et al., 2020) and hence exceeding the 2-year baseline of our programme.
@@
To avoid crowding, we removed objects that have a Gaia source brighter than mag closer than 1.2”, which corresponds to the FLAMES fibre size. We also explicitly excluded stars within 30” around the centres of the dense SMC clusters NGC 330 and NGC 346.
@@
In an attempt to select only massive stars, we made use of an evolutionary track of a star computed by Schootemeijer et al. (2019) with the Modules for Experiments in Stellar Astrophysics (MESA) stellar evolution code (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). We converted the physical parameters along the track to a -band magnitude and colours using bolometric corrections taken from the MIST webpage222http://waps.cfa.harvard.edu/MIST/model_grids.html; the bolometric corrections were retrieved by fitting a 5th order polynomial to MIST values for and . (Dotter, 2016; Choi et al., 2016). We adjusted the track on the color-magnitude diagram (CMD) by adopting a distance of 62 kpc (Graczyk et al., 2020), and assuming an average value for the reddening of mag and extinction of mag (Schootemeijer et al., 2021). We then only selected targets whose CMD positions lie above this track before it becomes a RSG (effective temperature kK; see Fig. 2). This resulted in a massive-star catalogue of 5576 stars subject to the criteria above. We also made use of a MESA track computed for by Schootemeijer et al. (2019) to divide the sample into stars with (born as O-type stars) and (born as B-type stars)333The threshold mass for O-type stars is typically taken as for the Galaxy (e.g. Martins et al., 2005). However, stars in the SMC are more compact and hot at a fixed mass (e.g. Georgy et al., 2013), such that this threshold is likely lower at low ..
@@
Our programme includes a total of eight FLAMES plate configurations, each with a field-of-view (FoV) of diameter 25’, although the instrument setup ensures visibility of targets only within a 20’ diameter (Fig. 1). For each FLAMES plate configuration, 130 giraffe fibres are available. We allocated 14 fibres for sky, leaving each field with 116 science targets. The only exception is field 8, for which 13 sky fibres and 117 science targets are available. This makes a total of science targets.
@@
First, the centre of a FLAMES pointing was chosen. The choice of field centre followed automatically by identifying the coordinate which encloses as many massive and bright stars as possible within a 25’ diameter, after removing stars which were already allocated in previous field allocations. Specifically, the pointings were selected by identifying the centre coordinates that result in the largest number of stars with (see above) and mag. This resulted in fields 1 – 8 shown in Fig. 1, which provide a good coverage of the massive-star content of the SMC. We note that while some of the fields overlap (e.g., 1 & 4), there is no overlap in the allocated stars within each field.
As a second step, for each plate configuration, targets within a 20’ diameter were allocated to the available fibres, starting from the brightest ones, while aiming at a homogeneous sampling across the -band. Specifically, we divided the magnitude range 10 – 16.5 mag into 15 bins, and aimed at a homogeneous coverage across these bins, resulting in stars per magnitude bin, per field. This could not always be achieved, given the rarity of bright stars, and the smaller parameter range of massive stars at lower magnitudes. Since the final fibre allocation depends on limitations related to the FLAMES fibre positioner, the remainder of the massive stars in each field in a diameter of 25’ were taken as backup targets, with their priority sorted by brightness.
The final fibre allocation was then performed using ESO’s Fibre Positioner Observation Support Software (FPOSS). The majority of our input targets made it to the final allocation, but the final sample includes a few dozens of backup targets. Sky fibres were allocated to 14 fibres in each field (13 for field 8) selected from concentric rings around the field centre where no known Gaia source is located. Finally, guide stars were selected following requirements in the ESO FLAMES manual for cycle P112.
The final sample of 929 targets is shown on a CMD in Fig. 2, along with a magnitude histogram. We also show the completeness fraction with respect to the underlying SMC Gaia catalogue as a function of -band magnitude. Out of the 929 stars, and based on the single-star tracks in Fig. 2, 323 have initial masses above track (”born as O-type stars”) and 606 are below it (”born as B-type stars”). Evidently, the sample reaches a completeness fraction of for the sub-sample, and for the sample.
@@ 3 Observations and data reduction
At the time of writing, nine out of 25 epochs were obtained during the first semester and processed in the framework of the BLOeM survey. The field centres, along with the MJD values of the epochs so far acquired, are provided in Table 2 of Appendix A.
The data reduction was performed with the giraffe pipeline v. 2.16.11 under the ESO CPL environment (v. 3.13.5). Each exposure was split into two sub-exposures for a robust removal of cosmics (Sect. 3.5). The data reduction itself consisted of four steps: bias and dark subtraction, flatfield correction, and wavelength calibration. All spectra were resampled by the ESO CPL pipeline to a constant wavelength step of 0.2Å (see also Sect. 3.2) and science spectra were then sky-subtracted and corrected for the barycentric motion. As a final step, we resampled the individual spectra to a common wavelength grid and co-added the spectra of individual targets to boost S/N in order to improve spectral typing. The spectra have a spectral resolving power of and cover the spectral range Å, with a median S/N of per pixel and epoch; details are provided below. We provide further details below on specific aspects of the data reduction process444The reduced data and co-added spectra will be made available via ESO phase 3 upon termination of propriety time; they are currently available on http://www.astro.tau.ac.il/~tshenar/DR3/; please contact T. Shenar or J. Bodensteiner for credentials. .
@@ 3.1 Temporal sampling
The temporal sampling of the epochs is not strictly defined a-priori, to allow for sufficient scheduling flexibility. A minimum separation between each epoch of 1 d is requested, and no larger than 20 d to ensure that all epochs would be acquired within a semester. The typical separations between the epochs used here, acquired during Sep. 2023 – Dec. 2023, are days to weeks (see Table 2), for a total time baseline of 50 – 70 d, depending on the field. The fact that the acquisition will take place across four semesters ensures that both short-scale and long-scale variability will be covered by the survey. Since multiplicity analysis is beyond the scope of this current paper, we refrain from a complete characterisation and Fourier mapping of the temporal sampling here.
@@ 3.2 Wavelength calibration
We paid particular attention to the quality of the wavelength solution. For FLAMES giraffe, a reference ThAr calibration frame in the LR02 setup is obtained at the end of an observing night by illuminating each fibre on a given plate with the light of a ThAr lamp. As a result, each fibre of each epoch has its own calibration ThAr spectrum which is used by the CPL pipeline to produce a 2D polynomial dispersion solution. The wavelength calibration solution was performed in two steps. First we used the instrument model provided by the pipeline static calibration v2.16.11 to compute a first-guess solution and run a first iteration of the giwavecalibration CPL recipe. We modified the standard options to allow for a large detection window of 20 pixels at first, for five iterations, then we progressively reduced it to 15 and ultimately 10 in the remaining 5 iterations. We decreased the rejection threshold from 1.2 to 3, allowing us to retain more lines and provide a first wavelength solution with a root-mean-square (rms) residual of 0.45-0.56 pixels. By comparing ThAr spectra of different nights we noticed a slight drift in the wavelength solution in various parts of the ThAr spectrum, with a higher stability in the centre of the wavelength range and a larger epoch-to-epoch variation at the blue and red part of the wavelength solution. We estimated the internal consistency to be no better than a few km s-1. To improve on this, we performed a second iteration by using the first solution as a new input guess-solution and re-iterating the giwavecalibration recipe. This yields a final dispersion solution, characterised with a rms residual in the range of 0.22 to 0.25 pixels.
Three epochs of field 1 have been observed with the SIMCAL lamp on at the beginning of the survey. The SIMCAL lamp yields a set of five ThAr spectra spread across the detector and acquired simultaneously with the science observations. This setup was discontinued because the glow of the strongest ThAr lines impacted the adjacent fibres signal, leaving a noticeable imprint of ThAr on nearby sky and weak objects. Yet, we noticed no difference in the quality of the wavelength solution with or without the SIMCAL lamp.
@@
Finally, we investigated the stability and consistency of the individual dispersion solutions across the BLOeM dataset. We specifically investigated the inter-epoch stability for given fibres as well as the intra-epoch consistency across all fibres of a given epoch. For the first one, which informs us about the temporal stability of the data, we cross-correlated the wavelength-calibrated ThAr spectra of a given fibre and field with that of the same fibre and field across all the epochs obtained so far. We found maximum peak-to-peak differences to be of 130 m s-1, with a standard deviation below 50 m s-1. For the latter, which informs us about the consistency of the dispersion solution across the BLOeM data sets, we cross-correlated the ThAr spectra of all fibres for a given epoch and field with one arbitrary ThAr spectrum (fibre #10) for the epoch under consideration. While the peak-to-peak and rms variations are slightly larger (750 and 150 m s-1 in the worse case), they remain well within the specifications of the instrument. For a few nights, no ThAr calibration frame could be obtained in the morning following the observations. In such cases, we use the frame closest in time, typically the one from the morning before. In eight cases, however, we had to resort to ThAr calibration frames taken 30 to 40 h apart of the science. Yet, no noticeable difference in the quality of the calibration could be found, again suggesting that temporal drifts are limited.
3.3 Sky-subtraction
@@
Sky spectra were obtained simultaneously with the science integration through a set of fibres allocated to empty patches of sky. These are dubbed SKY fibres and record all background signal, including the moon, nebular and sky emission spectra depending on the wavelength regime. Hence not all signal recorded by SKY fibres are from the ‘sky’ itself, but we nonetheless adopt the generic denomination here.
@@
As described earlier, we typically used 14 SKY fibres in each field and kept the location constant across all epochs of a given field. We visually inspected the SKY spectra from each field and each epoch and flag spectra that seemed to be significantly higher than the median of the epoch and field. These are possibly contaminated by faint objects and therefore not representative of the true background signal. Once a sky location has been flagged as contaminated in any of the epochs of a given field, it is rejected from all epochs so that the median sky is always computed with the same set of input locations. In this process, we rejected 4, 3, 1 and 2 sky fibres for fields 1, 2, 5 and 6 respectively. The sky correction was finally performed by subtracting the median spectra of the ‘valid’ SKY fibres at the corresponding field and epoch. The error sky spectrum was computed as the rms around the median sky at each pixel after masking sky pixels affected by cosmic rays through a iterative filtering. The error sky spectrum was added quadratically to the error science non-sky-subtracted spectrum produced by the pipeline in order to compute the error spectrum of the sky-subtracted science data.
@@
Of importance, any nebular component in the sky spectra is obtained at the position of the SKY fibres. In our adopted observational setup, these are located tens of arcsecs to several arcminutes away from any science spectra hence they do not reflect the local nebular conditions of any science targets. The process of taking the median across 10 to 14 sky positions ensures that any local nebulosity in the SKY fibres is averaged out. As a corollary, the sky-subtracted science spectra are not corrected for nebulosity and therefore retain their nebular component.
@@ 3.4 Normalisation
The sky-subtracted spectra of each sub-exposure were individually and automatically normalised using a designated Python script to remove the underlying continuum, which is a combination of the stellar continuum, reddening, and instrumental response. Overall, the non-normalised spectra are well-behaved and have a smooth, typically monotonically decreasing behaviour across the spectral range. The normalisation was performed by fitting a polynomial to automatically-identified continuum points along the spectrum. Following several independent attempts, we found that a polynomial of degree eight provided the best results in terms of robust normalisation; lower polynomial degrees resulted at times in underfit regions, while higher degrees introduced wavy patterns in the spectra which can impact subsequent analysis.
@@
3.5 Cosmic correction and combination of sub-exposures
To boost the S/N, the two sub-exposures of each epoch and star were combined into one exposure. Moreover, we utilised the availability of two sub-exposures for a robust cosmic correction using a designated Python script. The process is as follows. First, positive flux outliers are identified in each of the two sub-exposures via the condition , where is the minimum of the error spectra of both sub-exposures. Cosmics are identified as points flagged as such in one spectrum, but not in the other – this ensures that intrinsic emission features (e.g., disk features, wind features, nebular lines) are not removed via this process. In principle, cosmics may be present in the same pixels in both sub-exposures, but such cases occur so rarely that this does not warrant further consideration. The cosmics are then removed by replacing each pixel identified as cosmic in one sub-exposure with the flux value of the same pixel in the second sub-exposure. The flux of the remaining pixels is formed by quadratically adding the two sub-exposures. The S/N per pixel ranges between 20 – 300, with a median value between 70 – 120, depending on the field (Fig. 3).

### v2

@@ 2 Sample selection
The BLOeM sample (see Fig. 1) was selected from the third Gaia data release catalogue (Gaia DR3, Gaia Collaboration et al., 2023). The catalogue was retrieved from the Gaia database using a search radius of around the SMC centre ( = 00:52:38, -72:48:01; epoch J2000). To achieve S/N per pixel (0.2 Å spectral bin), only stars with mag were retrieved. Foreground objects were filtered via two constraints. First, the parallax was required to be consistent with zero111Gaia is not sensitive enough to measure non-vanishing parallaxes within errors at the SMC distance. within 5, that is . Second, the proper motions of the stars were required to be consistent within 15 with the SMC proper motion ( and ) following Yang et al. (2019) and Schootemeijer et al. (2021).
@@
We omitted the 12 known SMC Wolf-Rayet (WR) stars (Neugent et al., 2018) from the sample, because they have been previously monitored (Foellmi et al., 2003; Hainich et al., 2015; Shenar et al., 2016; Schootemeijer et al., 2024). Finally, we omitted potential red supergiants (RSGs) from the sample by imposing mag: due to the large radii of RSGs, RSG binaries have periods that exceed a few years (e.g. Patrick et al., 2019; Neugent et al., 2020) and hence exceed the two-year baseline of our programme.
@@
To avoid crowding, we removed objects that have a Gaia source brighter than mag closer than 1.2”, which corresponds to the FLAMES fibre size. We also explicitly excluded stars within 30” of the centres of the dense SMC clusters NGC 330 and NGC 346.
we made use of an evolutionary track of a star (see Sect 5.4) computed by Schootemeijer et al. (2019) and Hastings et al. (2021) with
@@
In an attempt to select only massive stars, we made use of an evolutionary track of a star (see Sect. 5.4) computed by Schootemeijer et al. (2019) and Hastings et al. (2021) with the Modules for Experiments in Stellar Astrophysics (MESA) stellar evolution code (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). We converted the physical parameters along the track to a -band magnitude and colours using bolometric corrections taken from the MIST webpage222http://waps.cfa.harvard.edu/MIST/model_grids.html; the bolometric corrections were retrieved by fitting a fifth-order polynomial to MIST values for and . (Dotter, 2016; Choi et al., 2016). We adjusted the track on the colour–magnitude diagram (CMD) by adopting a distance of 62 kpc (Graczyk et al., 2020), and assuming an average value for the reddening of mag and extinction of mag (Schootemeijer et al., 2021). We then only selected targets whose CMD positions lie above this track before becoming a RSG (effective temperature kK; see Fig. 2). This resulted in a massive-star catalogue of 5576 stars subject to the criteria above. We also made use of a MESA track computed for by Schootemeijer et al. (2019) to divide the sample into stars with (born as O-type stars) and (born as B-type stars)333The threshold mass for O-type stars is typically taken as for the Galaxy (e.g. Martins et al., 2005). However, stars in the SMC are more compact and hot at a fixed mass (e.g. Georgy et al., 2013), such that this threshold is likely lower at low .. While massive stars are typically born as OB-type on the main sequence, they can appear as OBAF blue/yellow supergiants after leaving the main sequence (in addition to Wolf-Rayet stars and GMK red supergiants, which were omitted from our survey, as described above). Hence, we can expect the spectral types of the sample stars to span the entire OBAF range.
@@
Our programme includes a total of eight FLAMES plate configurations, each with a field-of-view (FoV) of 25’ in diameter, although the instrument setup ensures visibility of targets only within a 20’ diameter (Fig. 1). For each FLAMES plate configuration, 130 giraffe fibres are available. We allocated 14 fibres for sky, leaving each field with 116 science targets. The only exception is field 8, for which 13 sky fibres and 117 science targets are available. This makes a total of science targets.
@@
First, the centre of a FLAMES pointing was chosen. The choice of field centre followed automatically by identifying the coordinate that encloses as many massive and bright stars as possible within a circle of 25’ in diameter centred on that coordinate, after removing stars which were already allocated in previous field allocations. Specifically, the pointings were selected by identifying the centre coordinates that result in the largest number of stars with (see above) and mag. This resulted in fields 1 – 8 shown in Fig. 1, which provide a good coverage of the massive-star content of the SMC. We note that while some of the fields overlap (e.g. 1 and 4), there is no overlap between the allocated stars within each field.
As a second step, for each plate configuration, targets within a circle of 20’ in diameter were allocated to the available fibres, starting from the brightest ones, while aiming to achieve a homogeneous sampling across the -band. Specifically, we divided the magnitude range 10 – 16.5 mag into 15 bins, and aimed to achieve homogeneous coverage across these bins, resulting in stars per magnitude bin, per field. This was not always possible, given the rarity of bright stars, and the smaller parameter range of massive stars at lower magnitudes. As the final fibre allocation depends on limitations related to the FLAMES fibre positioner, the remaining massive stars in each field within a circle of 25’ in diameter were taken as backup targets, with their priority sorted by brightness.
The final fibre allocation was then performed using ESO’s Fibre Positioner Observation Support Software (FPOSS). The majority of our input targets made it to the final allocation, but the final sample includes a few dozen backup targets. Sky fibres were allocated to 14 fibres in each field (13 for field 8) selected from concentric rings around the field centre where no known Gaia source is located. Finally, guide stars were selected following requirements in the ESO FLAMES manual for cycle P112.
The final sample of 929 targets is shown on a CMD in Fig. 2, along with a magnitude histogram. We also show the completeness fraction with respect to the underlying SMC Gaia catalogue as a function of -band magnitude. Of the 929 stars, and based on the single-star tracks in Fig. 2, 323 have initial masses of above (”born as O-type stars”) and 606 are below this mass (‘born as B-type stars’). Evidently, the sample reaches a completeness fraction of for the subsample, and for the sample.
@@ 3 Observations and data reduction
At the time of writing, 9 out of 25 epochs were obtained during the first semester and were processed in the framework of the BLOeM survey. The field centres, along with the MJD values of the epochs acquired so far, are provided in Table 2 of Appendix A.
The data reduction was performed with the giraffe pipeline v. 2.16.11 under the ESO CPL environment (v. 3.13.5). Each exposure was split into two subexposures for a robust removal of cosmic rays (cosmics, Sect. 3.5). The data reduction itself consisted of four steps: bias and dark subtraction, flatfield correction, and wavelength calibration. All spectra were resampled by the ESO CPL pipeline to a constant wavelength step of 0.2Å (see also Sect. 3.2) and science spectra were then sky-subtracted and corrected for the barycentric motion. As a final step, we resampled the individual spectra to a common wavelength grid and co-added the spectra of individual targets to boost S/N in order to improve spectral typing. The spectra have a spectral resolving power of and cover the spectral range Å, with a median S/N of per pixel and epoch; details are provided below. We provide further details below on specific aspects of the data reduction process444The reduced data and co-added spectra will be made available via ESO phase 3 upon termination of propriety time; they are currently available on http://www.astro.tau.ac.il/~tshenar/DR3/; please contact T. Shenar or J. Bodensteiner for credentials. .
@@ 3.1 Temporal sampling
The temporal sampling of the epochs is not strictly defined a priori in order to allow for sufficient scheduling flexibility. We insist on a minimum separation between each epoch of 1 d, and a maximum of 20 d to ensure that all epochs are acquired within a semester. The typical separations between the epochs used here, acquired during September 2023 through December 2023, are days to weeks (see Table 2), for a total time baseline of 50 – 70 d, depending on the field. The fact that the acquisition will take place across four semesters ensures that both short-scale and long-scale variability will be covered by the survey. As a multiplicity analysis is beyond the scope of the present paper, we refrain from a complete characterisation and Fourier mapping of the temporal sampling here.
@@ 3.2 Wavelength calibration
We paid particular attention to the quality of the wavelength solution. For FLAMES giraffe, a reference ThAr calibration frame in the LR02 setup is obtained at the end of an observing night by illuminating each fibre on a given plate with the light of a ThAr lamp. As a result, each fibre of each epoch has its own calibration ThAr spectrum, which is used by the CPL pipeline to produce a 2D polynomial dispersion solution. The wavelength calibration solution was performed in two steps. First we used the instrument model provided by the pipeline static calibration v2.16.11 to compute a first-guess solution and run a first iteration of the giwavecalibration CPL recipe. We modified the standard options to allow for a large detection window of 20 pixels at first, for five iterations, and then progressively reduced it to 15 and ultimately 10 in the remaining five iterations. We decreased the rejection threshold from 1.2 to 3, allowing us to retain a greater number of lines and provide a first wavelength solution with a root mean square (rms) residual of 0.45-0.56 pixels. By comparing ThAr spectra of different nights we noticed a slight drift in the wavelength solution in various parts of the ThAr spectrum, with a higher stability in the centre of the wavelength range and a larger epoch-to-epoch variation in the blue and red parts of the wavelength solution. We estimated the internal consistency to be no better than a few km s-1. To improve on this, we performed a second iteration using the first solution as a new input guess-solution and reiterating the giwavecalibration recipe. This yields a final dispersion solution, characterised with a rms residual in the range of 0.22 to 0.25 pixels.
Three epochs of field 1 were observed with the SIMCAL lamp on at the beginning of the survey. The SIMCAL lamp yields a set of five ThAr spectra spread across the detector and acquired simultaneously with the science observations. This setup was discontinued because the glow of the strongest ThAr lines impacted the signals of the adjacent fibres, leaving a noticeable imprint of ThAr on nearby sky and weak objects. Yet, we noticed no difference in the quality of the wavelength solution with or without the SIMCAL lamp.
@@
Finally, we investigated the stability and consistency of the individual dispersion solutions across the BLOeM dataset. We specifically investigated the inter-epoch stability for given fibres as well as the intra-epoch consistency across all fibres of a given epoch. For the first one, which informs us about the temporal stability of the data, we cross-correlated the wavelength-calibrated ThAr spectra of a given fibre and field with that of the same fibre and field across all the epochs obtained so far. We found maximum peak-to-peak differences to be of 130 m s-1, with a standard deviation of below 50 m s-1. For the latter, which informs us about the consistency of the dispersion solution across the BLOeM data sets, we cross-correlated the ThAr spectra of all fibres for a given epoch and field with one arbitrary ThAr spectrum (fibre #10) for the epoch under consideration. While the peak-to-peak and rms variations are slightly larger (750 and 150 m s-1 in the worse case), they remain well within the specifications of the instrument. For a few nights, no ThAr calibration frame could be obtained in the morning following the observations. In such cases, we use the frame closest in time, typically the one from the morning before. However, in eight cases we had to resort to ThAr calibration frames taken 30 to 40 h before of after observations of our targets. Nevertheless, no noticeable difference in the quality of the calibration could be found, again suggesting that temporal drifts are limited.
3.3 Sky subtraction
@@
Sky spectra were obtained simultaneously with the science integration through a set of fibres allocated to empty patches of sky. These are dubbed SKY fibres and record all background signal, including the moon and nebular and sky emission spectra depending on the wavelength regime. Hence, not all signal recorded by SKY fibres is from the ‘sky’ itself, but we nonetheless adopt the generic denomination here.
@@
As described earlier, we typically used 14 SKY fibres in each field and kept the location constant across all epochs of a given field. We visually inspected the SKY spectra from each field and each epoch and flagged spectra that seemed to be significantly higher than the median of the epoch and field. These are possibly contaminated by faint objects and therefore not representative of the true background signal. Once a sky location has been flagged as contaminated in any of the epochs of a given field, it is rejected from all epochs so that the median sky is always computed with the same set of input locations. In this process, we rejected 4, 3, 1, and 2 sky fibres for fields 1, 2, 5, and 6, respectively. The sky correction was finally performed by subtracting the median spectra of the ‘valid’ SKY fibres at the corresponding field and epoch. The error sky spectrum was computed as the rms around the median sky at each pixel after masking sky pixels affected by cosmic rays through a iterative filtering. The error sky spectrum was added quadratically to the error science non-sky-subtracted spectrum produced by the pipeline in order to compute the error spectrum of the sky-subtracted science data.
@@
Of importance, any nebular component in the sky spectra is obtained at the position of the SKY fibres. In our adopted observational setup, these are located tens of arcsecs to several arcminutes away from any science spectra and hence they do not reflect the local nebular conditions of any science targets. The process of taking the median across 10 to 14 sky positions ensures that any local nebulosity in the SKY fibres is averaged out. As a corollary, the sky-subtracted science spectra are not corrected for nebulosity and therefore retain their nebular component.
@@ 3.4 Normalisation
The sky-subtracted spectra of each subexposure were individually and automatically normalised using a designated Python script to remove the underlying continuum, which is a combination of the stellar continuum, reddening, and instrumental response. Overall, the non-normalised spectra are well-behaved and have a smooth, typically monotonically decreasing behaviour across the spectral range. The normalisation was performed by fitting a polynomial to automatically identified continuum points along the spectrum. Following several independent attempts, we found that a polynomial of degree eight provided the best results in terms of robust normalisation; lower polynomial degrees resulted at times in underfit regions, while higher degrees introduced wavy patterns in the spectra, which can impact subsequent analysis.
@@
3.5 Cosmic correction and combination of subexposures
To boost the S/N, the two subexposures of each epoch and star were combined into one exposure. Moreover, we used the availability of two subexposures for a robust cosmic correction using a designated Python script. The process is as follows. First, positive flux outliers are identified in each of the two subexposures via the condition , where is the minimum of the error spectra of both subexposures. Cosmics are identified as points flagged as such in one spectrum but not in the other – this ensures that intrinsic emission features (e.g. disc features, wind features, nebular lines) are not removed via this process. In principle, cosmics may be present in the same pixels in both subexposures, but such cases occur so rarely that this does not warrant further consideration. The cosmics are then removed by replacing each pixel identified as cosmic in one subexposure with the flux value of the same pixel in the second subexposure. The flux of the remaining pixels is formed by quadratically adding the two subexposures. The S/N per pixel ranges between 20 and 300, with a median value of between 70 and 120, depending on the field (Fig. 3).